Let T : R3 -> R3 be the linear operator
defined by T(x1, x2, x3) = (x1, x3, -2x2 - x3).
Let f(x) = - x³ + 2. Find the operator f(T).
1
Expert's answer
2020-11-18T19:07:37-0500

Given that "T:\R^3\rightarrow \R^3" is linear operator given by


"T(x_1,x_2,x_3)=(x_1,x_3,-2x_2-x_3)"

Consider the standard basis of "\R^3" which are "e_1,e_2,e_3" , thus


"T(e_1)=a_1e_1+a_2e_2+a_3e_3=(a_1,a_2,a_3)=(1,0,0)\\
T(e_2)=b_1e_1+b_2e_2+b_3e_3=(b_1,b_2,b_3)=(0,0,-2)\\
T(e_3)=c_1e_1+c_2e_2+c_3e_3=(c_1,c_2,c_3)=(0,1,-1)"

Thus, matrix of T with respect to standard basis is


"M=M(T)_v^v=\begin{bmatrix}
 1 &0&0\\
0&0&-2\\
0&1&-1
\end{bmatrix}"

Where "v=(e_1,e_2,e_3)" .


Now, polynomial of T is given "f(x)=-x^3+2" , thus operator


"f(T)=-T^3+3I\iff -M^3+2I_{3\times 3}"

Now,

"M^3=\left( \begin{array}{ccc} 1 & 0 & 0 \\ 0 & 2 & 2 \\ 0 & -1 & 3 \end{array} \right)\\
\implies -M^3=\left( \begin{array}{ccc} -1 & 0 & 0 \\ 0 &- 2 & -2 \\ 0 & 1 & -3 \end{array} \right)\\
\implies -M^3+2I_{3\times 3}=\left( \begin{array}{ccc} 1 & 0 & 0 \\ 0 & 0& -2 \\ 0 & 1 & -1\end{array} \right)\\"

Thus,


"f(T)=\left( \begin{array}{ccc} 1 & 0 & 0 \\ 0 & 0& -2 \\ 0 & 1 & -1\end{array} \right)\\"